Using the data below find the slope and intercept for the least squares        regression line. (SHOW ALL STEPS/WORK)

x

y

xy

x2

y2

16

109

1744

256

11881

25

122

3050

625

14884

39

143

5577

1521

20449

45

132

5940

2025

17424

49

199

9751

2401

39601

  64

185

11840

4096

34225

70

199

13930

4900

39601

29

130

3770

841

16900

57

175

9975

3249

30625

22

118

2596

484

13924

416

1512

68173

20398

239514
Answer #1

Solution:

X Y XY X^2 Y^2
16 109 1744 256 11881
25 122 3050 625 14884
39 143 5577 1521 20449
45 132 5940 2025 17424
49 199 9751 2401 39601
64 185 11840 4096 34225
70 199 13930 4900 39601
29 130 3770 841 16900
57 175 9975 3249 30625
22 118 2596 484 13924
416 1512 628992 173056 2286144

PREDICTED Y= a +bX b=rSD a=7 – 6X

n 11
sum(XY) 697165.00
sum(X) 832.00
sum(Y) 3024.00
sum(X^2) 193454.00
sum(Y^2) 2525658.00
Numerator 5152847.00
Denominator 5172948.48
r 0.9961
r square 0.9922
Xbar(mean) 75.6364
Ybar(mean) 274.9091
SD(X) 15.7233
SD(Y) 32.7804
b1 3.5889
b0 3.4571

Y-hat = b0 + bl(x)

The slope of regression line is b1 = 3.5889

The intercept of regression line is b0 = 3.4571